# Stanford CA 94305, USA

### v2

@@ 7 Guaranteed automatic integration library
Most of the uncertainty quantification methods in this article are about finding some value $\epsilon$ where, after sampling $f$ $n$ times, we have reason to believe that $|\hat{\mu}_{n}-\mu|\leqslant\epsilon$. The evidence behind that belief might be a certificate, a confidence interval or an asymptotic confidence interval. A complementary approach is to start with a target value of $\epsilon$ and look for a value of $n$ for which we will have reason to believe that $|\hat{\mu}_{n}-\mu|\leqslant\epsilon$. The Guaranteed Automatic Integration Library (GAIL) [82] allows us to choose $\epsilon$ and then sample sequentially until we are confident that $|\hat{\mu}_{n}-\mu|\leqslant\epsilon$ has been attained.
In GAIL, the user specifies $\epsilon$ and then the library is designed to return an interval $[a,b]$ of width no more than $2\epsilon$ with $a\leqslant\mu\leqslant b$. A probabilistic version, delivers an interval with $\Pr(a\leqslant\mu\leqslant b)\geqslant 1-\alpha$.
A first approach to GAIL using MC is in [44]. As noted in Section 3 such an approach requires some extra knowledge about $f({\bm{x}})$. The approach in [44] is to use a preliminary sample to get a probabilistic upper bound for $\sigma$. If $\Pr(\sigma\leqslant\hat{\sigma})\geqslant 1-\alpha_{1}$ for $\alpha_{1}<\alpha$ then we only need to find $a<b$ with $\Pr(a\leqslant{\mathbb{E}}(f({\bm{x}})\leqslant b))\geqslant 1-\alpha_{2}$ that holds whenever $\operatorname{Var}(f({\bm{x}}))\leqslant\hat{\sigma}^{2}$ and $\alpha_{1}+\alpha_{2}\leqslant\alpha$.
@@
A QMC version of GAIL is in [45] . It is based on the Sobol’ sequence and a Walsh decomposition of the integrand. See [23] for both of those. The Walsh decomposition expands $f$ into a sum of Walsh functions indexed by ${\bm{k}}\in\{0,1,2,\dots\}^{d}$ and then multiplied by coefficients $\hat{f}_{{\bm{k}}}$. The integration error can be bounded by a sum of $|\hat{f}_{{\bm{k}}}|$ over certain indices ${\bm{k}}$ that depend on which Sobol’ points are used. Given a sample from a Sobol’ sequence it is possible to estimate some of the Walsh function coefficients. They make an assumption about how Walsh coefficients decay as ${\bm{k}}$ moves farther from $\bm{0}$. That assumption places $f$ inside a cone and it lets them bound the sum of the remaining absolute coefficients given the estimated sum of absolute coefficients. They keep doubling the number of Sobol’ points and updating the error bound until the error bound is small enough.
@@ 8 New directions
A second challenge with UQ for MCMC is that their estimates typically have a bias. That bias may disappear exponentially fast with $n$ while remaining large in practice because $A\rho^{n}$ for $\rho$ just barely smaller than one may be large for the $n$ we can use. It is much easier to quantify uncertainty in unbiased estimates that can be independently replicated. There is some recent work on coupling strategies that remove the bias from MCMC. See [3] for a discussion of those methods. That allows independent replicates of unbiased MCMC methods to be used for UQ.
@@
By combining unbiased MCMC with a weakly CUD driving sequence, [26] are able to get independent unbiased replicates of MCMC estimates for the Gibbs sampler. That allows simple replication for MCMC with empirically better convergence than by plain MCMC. Much of the analysis extends beyond the Gibbs sampler, but the Gibbs sampler is smooth which can help it benefit from more evenly distributed inputs and their algorithm also uses a coupling strategy designed for the Gibbs sampler.
@@
For analytic functions on $[0,1]^{d}$, a median-of-means approach brings an error of $O(n^{-c\log_{2}(n)/d})$ for any $c<3\log(2)/\pi^{2}\approx 0.21$ [75]. This rate is called super-polynomial because it is better than $O(n^{-r})$ for any finite $r$. A median-of-means strategy adapts to possibly unknown level of smoothness in $f$. For $R\geqslant\log_{2}(n)$, it attains an RMSE of $O(n^{-\alpha-1/2+\epsilon})$ for any $\epsilon>0$ when $f$ has finite variation of order $\alpha$ (defined in [20]). This is proved in Theorem 2 of [72].
There are some other applications of the median of means method that attain a universal goodness property. Here are sketches of two of them; the reader should read them to get the full details. A median of means approach to RQMC by lattice rules has been developed by [36]. Lattice rules can be tuned to specific Hilbert space weights, but even if the weights are very badly chosen, a median of means estimate will do well. Goda and Krieg [35] choose a random prime number $p$ from the interval $[\lfloor n/2\rfloor+1,n]$. Then they choose a rank one lattice in $[0,1]^{d}$ with $p$ points using a generating vector chosen uniformly from $\{1,2,\dots,p-1\}^{d}$. From that rank one lattice, they estimate $\mu$. They repeat this $R=\Omega(\log(n))$ times independently and take $\hat{\mu}=\mathrm{median}(\hat{\mu}^{(1)},\dots,\hat{\mu}^{(R)})$. The resulting $\hat{\mu}$ is nearly optimal in any Korobov class of periodic functions with smoothness $\alpha>1/2$.
@@
The median of means method is most commonly used on $n$ IID random variables $Y_{i}$ with finite variance $\sigma^{2}$. The emphasis is very different from our use taking the median of $R$ independent estimates that each use $n$ very dependent values. While a CLT gives asymptotic confidence intervals for $\mu={\mathbb{E}}(Y_{i})$ the goal in a much median of means research is to get the desired coverage for finite $n$. That requires somewhat wider confidence interval than the CLT based ones and also is only available for $\alpha\geqslant\alpha_{\min}>0$ for a threshold $\alpha_{\min}$ that depends on some assumptions about the distribution of $Y_{i}$. See [18] and references therein. The methods use known $\sigma$ though an upper bound or upper confidence limit for $\sigma$ (as in GAIL) could be used.
@@
Other RQMC methods and other smoothness conditions can impose less stringent requirements on $R$ for a CLT. For example the digital nets of [31] scrambled as in [66] yield a CLT for $\hat{\mu}$ as $n\to\infty$ for fixed finite $R$ [57]. When UQ is the primary criterion with accuracy secondary, then those $t=0$ nets might be preferable to the ones of Sobol’ with $t>0$.
@@ 9 Conclusions
We have seen that uncertainty quantification for QMC estimates is subject to a complex mix of gaps and tradeoffs and surprises. The bracketing inequalities for smooth convex integrands are simple to use and very easy to understand. However they require strong assumptions and exhibit a severe dimension effect. The other methods with certificates also require strong assumptions and show a dimension effect. The methods based on NNLD and NPLD points use more involved constructions and more complex derivations.
@@ Acknowledgments
This work was supported by the National Science Foundation under grant DMS-2152780. I thank the people behind MCQMC 2024: Christiane Lemieux, Ben Feng, Nathan Kirk, Adam Kolkiewicz, Carla Daniels and Greg Preston for organizing such a delightful meeting. Most of the new work reported here was done in collaboration with Pierre L’Ecuyer, Bruno Tuffin, Marvin Nakayama, Michael Gnewuch, Peter Kritzer and Zexin Pan.

### v3

@@ 7 Guaranteed automatic integration library
Most of the uncertainty quantification methods in this article are about finding some value $\epsilon$ where, after sampling $f$ $n$ times, we have reason to believe that $|\hat{\mu}_{n}-\mu|\leqslant\epsilon$. The evidence behind that belief might be a certificate, a confidence interval or an asymptotic confidence interval. A complementary approach is to start with a target value of $\epsilon$ and look for a value of $n$ for which we will have reason to believe that $|\hat{\mu}_{n}-\mu|\leqslant\epsilon$. There are many results in the literature showing how $n$ must grow asymptotically as $\epsilon$ is reduced. The Guaranteed Automatic Integration Library (GAIL) [96] may be the unique one that provides a non-asymptotic solution for integration over $[0,1]^{d}$. The GAIL project continues as part of the qmcpy Python library [15].
In GAIL, the user specifies $\epsilon$ and then the library is designed to return an interval $[a,b]$ of width no more than $2\epsilon$ with $a\leqslant\mu\leqslant b$. In other words, GAIL provides a bracketing solution where the user can specify the width of the bracket instead of specifying $n$. A probabilistic version delivers an interval with $\Pr(a\leqslant\mu\leqslant b)\geqslant 1-\alpha$.
A first approach to GAIL using MC is in [50]. As noted in Section 3 such an approach requires some extra knowledge about $f({\bm{x}})$. The approach in [50] is to use a preliminary sample to get a probabilistic upper bound for $\sigma$. If $\Pr(\sigma\leqslant\hat{\sigma})\geqslant 1-\alpha_{1}$ for $\alpha_{1}<\alpha$ then we only need to find $a<b$ with $\Pr(a\leqslant{\mathbb{E}}(f({\bm{x}}))\leqslant b)\geqslant 1-\alpha_{2}$ that holds whenever $\operatorname{Var}(f({\bm{x}}))\leqslant\hat{\sigma}^{2}$ and $\alpha_{1}+\alpha_{2}\leqslant\alpha$.
@@
A QMC version of GAIL is in [51] . It is based on the Sobol’ sequence and a Walsh decomposition of the integrand. See [28] for both of those. The Walsh decomposition expands $f$ into a sum of Walsh functions indexed by ${\bm{k}}\in\{0,1,2,\dots\}^{d}$ and multiplied by coefficients $\hat{f}_{{\bm{k}}}$. The integration error can be bounded by a sum of $|\hat{f}_{{\bm{k}}}|$ over certain indices ${\bm{k}}$ that depend on which Sobol’ points are used. Given a sample from a Sobol’ sequence it is possible to estimate some of the Walsh function coefficients. They make an assumption about how Walsh coefficients decay as ${\bm{k}}$ moves farther from $\bm{0}$. Their assumption goes beyond the decay of Walsh coefficients noted by [24] and [102] for smooth $f$. That assumption places $f$ inside a cone and it lets them bound the sum of the remaining absolute coefficients given the estimated sum of absolute coefficients. They keep doubling the number of Sobol’ points and updating the error bound until the error bound is small enough.
A counterpart for rank one lattices, using a Fourier decomposition is in [54]. A discussion of relative error is studied in [95].
@@ 8 New directions
A second challenge with UQ for MCMC is that their estimates typically have a bias. That bias may disappear exponentially fast with $n$ while remaining large in practice because $A\rho^{n}$ for $\rho$ just barely smaller than one may be large for the $n$ we can use. It is much easier to quantify uncertainty in unbiased estimates that can be independently replicated. Coupling from the past (CFTP) [88] can generate unbiased estimates in some MCMC settings, but it only works well in very restrictive settings. There is some recent work on more generally applicable coupling strategies that remove the bias from MCMC. See [3] for a discussion of those methods. That allows independent replicates of unbiased MCMC methods to be used for UQ.
@@
By combining unbiased MCMC with a weakly CUD driving sequence, [31] are able to get independent unbiased replicates of MCMC estimates for the Gibbs sampler. That allows simple replication for MCMC with empirically better convergence than by plain MCMC. Much of the analysis extends beyond the Gibbs sampler, but the Gibbs sampler is smooth which can help it benefit from more evenly distributed inputs and their algorithm also uses a coupling strategy designed for the Gibbs sampler. When CFTP is applicable, it can be used with RQMC [62].
@@
For analytic functions on $[0,1]^{d}$, a median-of-means approach brings an error of $O(n^{-c\log_{2}(n)/d})$ for any $c<3\log(2)/\pi^{2}\approx 0.21$ [86]. This rate is called super-polynomial because it is better than $O(n^{-r})$ for any finite $r$. A median-of-means strategy adapts to a possibly unknown level of smoothness in $f$. For $R\geqslant\log_{2}(n)$, it attains an RMSE of $O(n^{-\alpha-1/2+\epsilon})$ for any $\epsilon>0$ when $f$ has finite variation of order $\alpha$ (defined in [25]). This is proved in Theorem 2 of [83].
There are some other applications of the median of means method that attain a universal goodness property. Here are sketches of two of them; the reader should read them to get the full details. A median of means approach to RQMC by lattice rules has been developed by [42]. Lattice rules can be tuned to specific Hilbert space weights. The median rules in [42] can be constructed without specifying those weights and yet they attain nearly the optimal worst-case error rate for any reasonable choice of weights and smoothness. Goda and Krieg [41] choose a random prime number $p$ from the interval $[\lfloor n/2\rfloor+1,n]$. Then they choose a rank one lattice in $[0,1]^{d}$ with $p$ points using a generating vector chosen uniformly from $\{1,2,\dots,p-1\}^{d}$. From that rank one lattice, they estimate $\mu$. They repeat this $R=\Omega(\log(n))$ times independently and take $\hat{\mu}=\mathrm{median}(\hat{\mu}^{(1)},\dots,\hat{\mu}^{(R)})$. The resulting $\hat{\mu}$ is nearly optimal in any Korobov class of periodic functions with smoothness $\alpha>1/2$.
@@
The median of means method is most commonly used on $n$ IID random variables $Y_{i}$ with finite variance $\sigma^{2}$. The emphasis is very different from our use taking the median of $R$ independent estimates that each use $n$ very dependent values. While a CLT gives asymptotic confidence intervals for $\mu={\mathbb{E}}(Y_{i})$ the goal in much median of means research is to get the desired coverage for finite $n$. That requires somewhat wider confidence interval than the CLT based ones and also is only available for $\alpha\geqslant\alpha_{\min}>0$ for a threshold $\alpha_{\min}$ that depends on some assumptions about the distribution of $Y_{i}$. See [22] and references therein. The methods use known $\sigma$ though an upper bound or upper confidence limit for $\sigma$ (as in GAIL) could be used.
@@
Other RQMC methods and other smoothness conditions can impose less stringent requirements on $R$ for a CLT. For example the digital nets of [36] scrambled as in [76] are known to yield a CLT for $\hat{\mu}$ as $n\to\infty$ for fixed finite $R$ [67]. When UQ is the primary criterion with accuracy secondary, then those $t=0$ nets might be preferable to the ones of Sobol’ with $t>0$ where no CLT has been proven.
@@ 9 Conclusions
We have seen that uncertainty quantification for QMC estimates is subject to a complex mix of gaps and tradeoffs and surprises. The bracketing inequalities for smooth convex integrands are simple to use and very easy to understand. However they require strong assumptions, exhibit a severe dimension effect and can greatly overestimate the size of the error. The other methods with certificates also require strong assumptions and show a dimension effect. The methods based on NNLD and NPLD points use more involved constructions and more complex derivations.
@@ Acknowledgments
This work was supported by the National Science Foundation under grant DMS-2152780. I thank the people behind MCQMC 2024: Christiane Lemieux, Ben Feng, Nathan Kirk, Adam Kolkiewicz, Carla Daniels and Greg Preston for organizing such a delightful meeting. Most of the new work reported here was done in collaboration with Pierre L’Ecuyer, Bruno Tuffin, Marvin Nakayama, Michael Gnewuch, Peter Kritzer and Zexin Pan. I received helpful comments on this document from Peter Kritzer, Marvin Nakayama, Pierre L’Ecuyer and two anonymous reviewers.